# Joint distribution of Hecke eigenforms on $\mathbbm{H}^{3}$

### v1

@@ 3.3. Bounds on L-factors
The expressions obtained are similar to the $\mathrm{GL}_{2}(\mathbbm{Q})$ case of Huang [HHL24], with $\Gamma_{\mathbbm{C}}(s)$ replacing $\Gamma_{\mathbbm{R}}(s)=\pi^{-s/2}\Gamma(\tfrac{s}{2})$. As a result, after applying Lemma 2.6, it is possible to get similar estimates to [HHL24], with different polynomial terms. We report the explicit expressions for completeness.
@@ 3.4. Final Estimates
The other terms in (46) can be treated in a similar way. The main input comes from the study of the function $Q_{1}(t_{j},t_{f},t_{g},t_{k})$. From (58), we get that if $Q_{1}$ is large enough, the exponential contribution will dominate the polynomial terms $P(t_{j},t_{f},t_{g},t_{k})$ coming from the convexity bounds available for the central values of the $L$-functions considered, (see [IK04]) and the logarithmic factors coming from the symmetric square $L$-functions (this is equivalent to the non-existence of Siegel zeros, and this is proven for symmetric squares of $\mathrm{GL}(2)$-forms in [HR95], conditionally on an hypothesis proven in [Ban97]). For this reason, the sum in (63) can be truncated to $t_{j}<2t_{f}+t_{k}+t_{g}$. In fact, in the other case we have $Q_{1}>2t_{f}+t_{k}+t_{j}$ and it grows linearly in $t_{j}$, which results in an exponential decay in $t_{j}$ for the considered terms, from (58). This implies that the contribution of the second range is negligible or comparable to the first range unconditionally.
@@
The methods of this paper generalize directly to the general case of number fields $F$ of class number one. The contribution given by the different infinite places in the analogue of Section 3.3 can be treated separately. The estimates finally reduce to $r$ copies of the real case as in [HHL24] and $s$ copies of the complex case treated in the present article, for $r$ and $s$ respectively the number of real and complex embeddings of $F$. The expected result is a bound on the error term of the form

### v2

@@ 3.3. Bounds on L-factors
The expressions obtained are similar to the $\mathrm{GL}_{2}(\mathbbm{Q})$ case of Hua-Huang-Li [hua_joint_2024], with $\Gamma_{\mathbbm{C}}(s)$ replacing $\Gamma_{\mathbbm{R}}(s)=\pi^{-s/2}\Gamma(\tfrac{s}{2})$. As a result, after applying Lemma 2.6, it is possible to get similar estimates to [hua_joint_2024], with different polynomial terms. We report the explicit expressions for completeness.
@@ 3.4. Final Estimates
The other terms in (44) can be treated in a similar way. The main input comes from the study of the function $Q_{1}(t_{j},t_{f},t_{g},t_{k})$. From (56), we get that if $Q_{1}$ is large enough, the exponential contribution will dominate the polynomial terms $P(t_{j},t_{f},t_{g},t_{k})$ coming from the convexity bounds available for the central values of the $L$-functions considered, (see [ik]) and the logarithmic factors coming from the symmetric square $L$-functions (this is equivalent to the non-existence of Siegel zeros, and this is proven for symmetric squares of $\mathrm{GL}(2)$-forms in [hoffstein_siegel_1995], conditionally on an hypothesis proven in [banks_twisted_1997]). For this reason, the sum in (61) can be truncated to $t_{j}<2t_{f}+t_{k}+t_{g}$. In fact, in the other case we have $Q_{1}>2t_{f}+t_{k}+t_{j}$ and it grows linearly in $t_{j}$, which results in an exponential decay in $t_{j}$ for the considered terms, from (56). This implies that the contribution of the second range is negligible or comparable to the first range unconditionally.
@@
The methods of this paper generalize directly to the general case of number fields $F$ of class number one. The contribution given by the different infinite places in the analogue of Section 3.3 can be treated separately. The estimates finally reduce to $r$ copies of the real case as in [hua_joint_2024] and $s$ copies of the complex case treated in the present article, for $r$ and $s$ respectively the number of real and complex embeddings of $F$. The expected result is a bound on the error term of the form
@@
Data availability statement. Data sharing is not applicable — the paper describes entirely theoretical research.
Conflict of interest statement The authors declare that they have no conflicts of interest.
Acknowledgments L. M. is Co-Funded by the European Union’s Horizon Europe research and innovation programme under the Marie Sklodowska-Curie Grant Agreement No 101126554. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them.